If 5 moles of N2 reacted with 12 moles of O2 according to 0 N2 + 2 O2 + NO2 how many moles NO2 are produced


1
Expert's answer
2021-05-20T07:51:49-0400

Solution:

Balanced chemical equation:

N2 + 2O→ 2NO2

According to the equation above: 2×n(N2) = n(O2) = n(NO2)


Moles of each reactant:

n(N2) = 5 mol

n(O2) = 12 mol


Choose one reactant and determine how many moles of the other reactant are necessary to completely react with it. Let's choose N2:

n(O2) = 2 × n(N2) = 2 × 5 mol = 10 mol


The calculation above means that we need 10 mol of O2 to completely react with N2.

We have 12 mol of O2 and therefore more than enough oxygen gas.

Thus oxygen gas (O2) is in excess and nitrogen gas (N2) must be the limiting reagent.


Therefore,

n(NO2) = 2 × n(N2) = 2 × 5 mol = 10 mol

Moles of NO2 = 10 mol


Answer: 10 mol of NO2 are produced.